# This is a preprint submitted to: Computer Methods in Applied Mechanics and Engineering

### v1

@@ Abstract
In this work, an extension of the parametric model embedding (PME) approach is presented, aiming to achieve more effective design-space dimensionality reduction for shape optimization in vehicle design. PME, rooted in principal component analysis (PCA), not only identifies a reduced set of critical modes but also re-parameterizes the original design space—enabling direct and interpretable manipulations of shape modifications within the reduced space. Alongside the “physics-informed” version (PI-PME), which enriches geometry with low-fidelity distributed and lumped physical quantities, a “physics-driven” variant (PD-PME) is introduced that focuses exclusively on physical parameters. Both formulations employ PCA to capture the principal modes of variability yet differ in their balance between geometric and physical information, through the ad-hoc definition of a weighted inner product. Through test cases involving the RAE-2822 airfoil, a bio-inspired underwater glider, a naval propeller, and the DTMB-5415 destroyer-type vessel, it is shown how the resulting frameworks provide a first-level assessment of design variability, offer interpretability regarding which original variables most strongly affect performance, and efficiently bridge geometric and physical parameters. Furthermore, lumped physical parameters can serve as a low-fidelity foundation for multi-fidelity optimization, directly leveraging the linear re-parameterization to drive the reduced design variables. Meanwhile, distributed physical parameters enable the construction of machine-learning-based reduced-order models to infer integral quantities of interest. By allowing the user to embed these insights early in the design process, PI-PME and PD-PME facilitate more robust, cost-effective exploration, paving the way for subsequent high-fidelity optimization.
@@ 1 Introduction
Design-space dimensionality reduction methods for shape optimization [2] have been extensively explored to address these challenges, simplifying high-dimensional design spaces without significantly compromising predictive accuracy. Linear methods, such as principal component analysis (PCA) [3, 4, 5, 6] and singular value decomposition (SVD) [7, 8, 9] derived from proper orthogonal decomposition (POD) [10, 11, 12, 13, 14, 15] and Karhunen-Loève expansion (KLE) [16, 17, 18], have been widely adopted for reducing the dimensionality of design space by capturing the dominant modes of geometric variance. However, traditional PCA-based approaches primarily utilize geometric data, potentially neglecting critical physical phenomena influencing design performance. To enhance predictive capabilities, physics-informed PCA methods have recently emerged [19, 20, 21, 22, 23], augmenting geometric data matrices with physical simulation outputs, such as lumped parameters (e.g., efficiency or drag coefficients) or distributed physical quantities (e.g., pressure or velocity fields). Furthermore, [24, 25, 26, 27, 28] introduced geometric moments as cost-effective, physics-related descriptors that substitute direct simulation data, further enriching the PCA model without the computational cost of numerical simulations. These enriched representations allow dimensionality reduction frameworks to capture not only geometric but also essential physical variability, significantly improving design relevance and effectiveness.
Among these methods, parametric model embedding (PME) [29] is a promising approach for shape parameterization and re-parameterization [30]. Rooted in PCA, PME reduces the dimensionality of geometric data, capturing only the essential features while preserving key parametric relationships among design variables. This feature makes PME particularly suitable for industrial applications and vehicle design, where compatibility with existing CAD-based modeling frameworks is essential. By providing an explicit mapping between the reduced and original design variables, PME addresses the so-called pre-image problem [31] that limits traditional dimensionality reduction methods. However, PME focuses primarily on geometric variations, potentially overlooking critical physical phenomena—such as lift, drag, and viscous effects—that are integral to the performance of vehicles operating in fluid environments.
@@
To overcome this limitation, an extended framework called physics-informed parametric model embedding (PI-PME) is here introduced. This approach enriches the data matrix with physical information, including lumped parameters (e.g., force coefficients or efficiencies) or distributed quantities (e.g., surface pressure fields), derived from low-fidelity simulations. Although this enrichment entails an upfront computational cost, the resulting latent space captures both geometric and physical variability, enabling a more informed and reliable assessment of performance trade-offs. By integrating physical insights at the earliest stages of design exploration, PI-PME provides a foundation that can guide subsequent high-fidelity simulations and optimizations.
@@
Building on the same principle, a physics-driven parametric model embedding (PD-PME) is also proposed, which goes one step further by focusing solely on physical parameters and effectively excluding geometric variance. Although this approach removes explicit shape information from the dimensionality reduction, it can be advantageous in scenarios where the primary objective is to characterize or optimize physical behaviors, and where the link between geometry and physics is well captured through simplified models. This exclusive emphasis on physics can also serve as a natural starting point for multi-fidelity optimization or for data-driven reduced-order modeling, where lumped or distributed physical data are used to predict key performance indicators at reduced computational expense.
@@
The effectiveness of PI-PME and PD-PME becomes particularly apparent in real-world applications such as bio-inspired underwater gliders, naval propellers, or classical airfoils (e.g., the RAE-2822). In these cases, fluid-dynamic effects strongly impact performance, and omitting them from the dimensionality reduction step may lead to suboptimal designs and increased resource expenditure. By embedding physical information—whether partially, as in PI-PME, or exclusively, as in PD-PME—into the reduced space, it becomes possible to filter the vast design domain more intelligently, ensuring that only the most relevant configurations are explored. This synergy between geometric and physical factors not only improves the quality of early design-stage decisions but also streamlines the path to final optimization, cutting down both computational time and overall development costs. It may be noted that PME’s flexible structure also makes it directly applicable to structural optimization. Instead of relying on aerodynamic or hydrodynamic simulations, one may incorporate structural properties (e.g., mass moments of inertia, stiffness distributions, or stress fields) into the PME framework, enabling integrated assessments of mechanical performance [20, 32].
@@
The following sections illustrate the mathematical underpinnings of PME, PI-PME, and PD-PME, including detailed discussions of how PCA is adapted to incorporate physical data, and how these frameworks can be applied to a range of high-dimensional vehicle design problems. Through examples, the benefits of early-stage integration of physical insights will be demonstrated, showing improved design robustness, a decrease in exhaustive computational campaigns, and the potential for more agile innovation cycles in modern vehicle development. It may finally be noted that, while this work focuses on vehicle applications, the proposed method is broadly applicable to general shape optimization problems involving functional surfaces influenced by physical constraints.
@@ 2 Design-space dimensionality reduction methods
In shape optimization problems aimed at minimizing a function $f({\mathbf{u}})$, epistemic uncertainty manifests as limited prior knowledge about where the global optimum lies within the design space. While it is assumed that such an optimum exists, its position is not known a priori, thus requiring an extensive exploration of the domain. To formalize this, the design variable vector ${\mathbf{u}}$ may be considered as belonging to a stochastic domain $\mathcal{U}$, with an associated probability density function $p({\mathbf{u}})$ that encodes the incomplete understanding of the feasible region. This uncertainty propagates through the modeling process, turning both geometric and physical parameters into random fields.
Dimensionality reduction methods—such as POD, KLE, SVD, and PCA—offer a systematic way to handle this high-dimensional, uncertain landscape. By isolating the principal modes of variation, these techniques distill large datasets and complex models into reduced representations that capture the key features most relevant to the optimization. In so doing, they mitigate the computational burden of exploring vast design spaces while preserving critical information about geometry and performance. As will be discussed in the following sections, integrating physical insights into these dimensionality reduction frameworks, for design-space assessment and subsequent shape optimization, provides a more robust path toward identifying optimal solutions under uncertainty.
@@ 2.1 Parametric model embedding
PME [29] is a design-space dimensionality reduction method that extends the standard PCA approach by incorporating both shape deformations and design variables into a generalized feature space. Specifically, PME applies PCA to an augmented matrix that includes the discretized shape deformation vector ${\mathbf{d}}$ and the original design variables ${\mathbf{u}}$. This extension allows PME to directly map the reduced design space back to the original design variables without the need for reparameterization, which is typically required in standard PCA approaches. As a result, PME offers a more robust and practical method for maintaining the integrity of the original design features while facilitating effective shape optimization.
@@
where ${\boldsymbol{\delta}}({\boldsymbol{\xi}},{\mathbf{u}})\in\mathbb{R}^{n}$ is the resulting shape modification vector, defined by arbitrary shape parameterization or modification method (e.g., CAD parameterization, Bezier surfaces, free-form deformation, NURBS, etc.), and ${\mathbf{u}}\in\mathcal{U}\subset\mathbb{R}^{M}$ is the design-variable vector of dimension $M$.
@@
$\mathbf{D}=\left[\begin{array}[]{ccc}\hat{{\mathbf{d}}}_{1}^{(1)}&\dots&\hat{{\mathbf{d}}}_{1}^{(S)}\\ \vdots&\vdots&\vdots\\ \hat{{\mathbf{d}}}_{n}^{(1)}&\dots&\hat{{\mathbf{d}}}_{n}^{(S)}\\ \end{array}\right]=\left[\begin{array}[]{ccc}\hat{d}_{1,\xi_{1}}({\mathbf{u}}_{1})&&\hat{d}_{1,\xi_{1}}({\mathbf{u}}_{S})\\ \vdots&&\vdots\\ \hat{d}_{L,\xi_{1}}({\mathbf{u}}_{1})&&\hat{d}_{L,\xi_{1}}({\mathbf{u}}_{S})\\ \vdots&\dots&\vdots\\ \hat{d}_{1,\xi_{n}}({\mathbf{u}}_{1})&&\hat{d}_{1,\xi_{n}}({\mathbf{u}}_{S})\\ \vdots&&\vdots\\ \hat{d}_{L,\xi_{n}}({\mathbf{u}}_{1})&&\hat{d}_{L,\xi_{n}}({\mathbf{u}}_{S})\\ \end{array}\right]$ (2)
@@
$\mathbf{P}=\left[\begin{array}[]{c}\mathbf{D}\\ \mathbf{U}\end{array}\right]\qquad\mathrm{with}\qquad\mathbf{U}=\left[\begin{array}[]{ccc}\hat{{\mathbf{u}}}^{(1)}&\dots&\hat{{\mathbf{u}}}^{(1)}\\ \end{array}\right]=\left[\begin{array}[]{ccc}\hat{u}_{1,1}&&\hat{u}_{1,S}\\ \vdots&\cdots&\vdots\\ \hat{u}_{M,1}&&\hat{u}_{M,S}\\ \end{array}\right]$ (3)
@@
$\mathbf{G}=\left[\begin{array}[]{cc}\mathbf{G}_{\mathbf{d}}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}\\ \end{array}\right]\qquad\mathrm{and}\qquad{\mathbf{Z}}=\left[{{\mathbf{z}}}_{1}\,\,\,\dots\,\,\,{{\mathbf{z}}}_{S}\right]\qquad\mathrm{with}\qquad{{\mathbf{z}}}_{k}=\left[\begin{array}[]{c}\mathbf{q}_{k}\\ {\mathbf{v}}_{k}\end{array}\right]$ (6)
Here, $\mathbf{q}_{k}$ and ${\mathbf{v}}_{k}$ represent the eigenvector components associated to the shape modification ${\mathbf{d}}$ and design variable ${\mathbf{u}}$ vectors, respectively. The matrix $\mathbf{G}_{\mathbf{d}}=\mathrm{diag}\left(\mathbf{G}_{1},\dots,\mathbf{G}_{n}\right)$ is block diagonal and has dimensionality $\left[nL\times nL\right]$, with each $\left[L\times L\right]$ $k$-th block being a diagonal matrix itself
@@
$\mathbf{Z}^{\star}=\mathbf{Z}\mathbf{\Gamma}^{-1},\quad\text{where }\mathbf{\Gamma}=\mathrm{diag}\left(\left[\sqrt{\mathbf{z}_{1}^{\mathsf{T}}\mathbf{G}\mathbf{W}\mathbf{z}_{1}},\dots,\sqrt{\mathbf{z}_{nL}^{\mathsf{T}}\mathbf{G}\mathbf{W}\mathbf{z}_{nL}}\right]\right).$ (10)
@@
Finally, the solutions $\lambda_{k}$ and the corresponding normalized eigenvectors $\mathbf{v}_{k}^{\star}$ (columns component of $\mathbf{Z}^{\star}$ associated to the original design variables, see Eq. 6) are used to construct the reduced dimensionality representation of the original parameterization by means of the $N$ reduced design variables $\mathbf{x}=[x_{1},\dots,x_{N}]^{\mathsf{T}}$. Defining the desired confidence level $l$, with $0<l\leq 1$, the number of reduced design variables $N$ is chosen such that
@@
To ensure that all the samples in $\mathbf{P}$ can be reconstructed through the reduced-order representation of the original design space, the reduced design variables ${\mathbf{x}}$ can be bounded by identifying the minimum and maximum values attained by each component $\Theta_{jk}$ of the projection coefficients $\boldsymbol{\theta}_{j}$, as follows:

### v2

# This is a preprint submitted to: Engineering with Computers

@@ Abstract
Design-space dimensionality reduction is essential to mitigate the cost of high-fidelity simulation-based optimization, especially when dealing with high-dimensional geometric parameterizations. Traditional linear techniques, such as principal component analysis, are widely used but often neglect the physical response of the system and lack invertibility to the design space, i.e., the ability to reconstruct the original design parameters from a reduced representation. This work introduces two physics-aware extensions of the parametric model embedding (PME) framework, aimed at generating reduced representations that incorporate physical information while maintaining analytical backmapping. The first, physics-informed PME (PI-PME), combines geometric and physical variability; the second, physics-driven PME (PD-PME), relies solely on physical responses. The proposed methods enable the construction of interpretable and physically relevant reduced spaces that can be used for design-space exploration, surrogate modeling, and optimization. The approach is demonstrated on multiple engineering configurations, including airfoils, propellers, gliders, and hulls, showing its ability to capture performance-relevant directions and preserve parametric consistency. The methodology is offline and non-intrusive, compatible with low-fidelity simulations, and requires only a modest number of samples to ensure variance convergence
@@ 1 Introduction
Design-space dimensionality reduction methods for shape optimization [2] have been extensively explored to address these challenges, simplifying high-dimensional design spaces without significantly compromising predictive accuracy and can be broadly categorized into offline and online approaches. Offline or upfront methods (e.g., principal component analysis and its variants [3, 4]) aim to construct a reduced representation before the actual optimization or design exploration process, relying on a fixed dataset of samples. Conversely, online or adaptive techniques (e.g., active subspace methods [5, 6]) seek to refine the reduced space on-the-fly, as new data becomes available. While online approaches may offer enhanced flexibility, they often require substantial computational overhead and complex integration with the optimization workflow. In contrast, offline methods are appealing for their simplicity, interpretability, and suitability for use with high-fidelity simulations, provided that the initial sampling ensures sufficient coverage of the design space. Among them, linear methods, such as principal component analysis (PCA) [3, 7, 8, 9] and singular value decomposition (SVD) [10, 11, 12] derived from proper orthogonal decomposition (POD) [13, 14, 15, 16, 17, 18] and Karhunen-Loève expansion (KLE) [4, 19, 20], have been widely adopted for reducing the dimensionality of design space by capturing the dominant modes of geometric variance. However, traditional PCA-based approaches primarily utilize geometric data, potentially neglecting critical physical phenomena influencing design performance. To enhance predictive capabilities, physics-informed PCA methods have recently emerged [21, 22, 23, 24, 25], augmenting geometric data matrices with physical simulation outputs, such as lumped parameters (e.g., efficiency or drag coefficients) or distributed physical quantities (e.g., pressure or velocity fields). Furthermore, [26, 27, 28, 29, 30] introduced geometric moments as cost-effective, physics-related descriptors that substitute direct simulation data, further enriching the PCA model without the computational cost of numerical simulations. These enriched representations allow dimensionality reduction frameworks to capture not only geometric but also essential physical variability, significantly improving design relevance and effectiveness.
Despite their widespread use, conventional offline dimensionality reduction techniques such as PCA and SVD present critical limitations in design applications. First, SVD-based approaches, which are often employed in POD and KLE formulations, do not support weighting of the inputs due to their inherent algebraic structure. This prevents users from tuning the relative importance of geometric features or physical observables during the reduction. In contrast, PCA—formulated as an eigenvalue problem of a covariance matrix—can accommodate weighted formulations, enabling generalized PCA variants and physics-informed extensions that incorporate additional information beyond shape geometry. However, these approaches still lack a crucial capability: the ability to reconstruct a feasible set of design parameters from a point in the reduced space. This issue, commonly referred to as the pre-image problem [31], severely limits their applicability in optimization workflows or CAD-integrated environments where direct parametric interpretation and reversibility are required.
To overcome the latter limitation, the parametric model embedding (PME) technique [32] was introduced. PME embeds the original parametric variables explicitly into the reduction process and re-parameterizes [33] the design space by constructing an orthonormal basis that supports both weighted formulations and analytical backmapping, thereby bridging the gap between statistical reduction and design interpretability.
Table 1: Comparison of dimensionality reduction methods in terms of data types handled, use of design variables, availability of weight control, and possibility of backmapping to the original parametric space.
Method Geometry Physics Variables Weights Backmapping Notes
SVD Standard decomposition of geometry-only snapshots.
PCA Can weight geometric features.
PI-PCA Can also incorporate physics thanks to weights.
PME Embeds geometry and variables.
PI-PME Combines geometry and physics with variables embedding.
PD-PME Fully physics-driven embedding.
Nevertheless, the original PME was designed to address the pre-image problem only, without accounting for physical integral quantities or distributed fields, which are essential to characterize performance in fluid-dynamic applications. The present work extends PME to incorporate such physical information, exploring two complementary linear offline strategies: physics-informed PME (PI-PME), which embeds physical outputs alongside geometry and design variables, and physics-driven PME (PD-PME), which relies solely on physical features to define the embedding.
@@
PI-PME enriches the data matrix with physical information, including lumped parameters (e.g., force coefficients or efficiencies) or distributed quantities (e.g., surface pressure fields), derived from low-fidelity simulations. Although this enrichment entails an upfront computational cost, the resulting latent space captures both geometric and physical variability, enabling a more informed and reliable assessment of performance trade-offs. By integrating physical insights at the earliest stages of design exploration, PI-PME provides a foundation that can guide subsequent high-fidelity simulations and optimizations.
@@
Building on the same principle, a PD-PME is also proposed, which goes one step further by focusing solely on physical parameters and effectively excluding geometric variance. Although this approach removes explicit shape information from the dimensionality reduction, it can be advantageous in scenarios where the primary objective is to characterize or optimize physical behaviors, and where the link between geometry and physics is well captured through simplified models. This exclusive emphasis on physics can also serve as a natural starting point for multi-fidelity optimization or for data-driven reduced-order modeling, where lumped or distributed physical data are used to predict key performance indicators at reduced computational expense.
Table 1 summarizes the distinguishing features of the considered dimensionality reduction methods with respect to four key aspects: (i) the type of data handled (geometry, physics, or both), (ii) the inclusion of design variables in the embedding, (iii) the availability of weight control, and (iv) the capability to perform backmapping to the original parametric space. Classical methods such as SVD and PCA rely solely on geometric information and, while PCA can incorporate weights to prioritize specific features, neither provides a mechanism for recovering design parameters from reduced coordinates. Physics-informed variants of PCA extend this formulation by integrating physical observables—either lumped or distributed—but still lack the ability to reconstruct input designs. In contrast, the PME framework embeds design variables directly into the reduction process, enabling analytical backmapping and supporting weighted formulations. Its extensions, PI-PME and PD-PME, further enrich the embedding by incorporating physical quantities. While PI-PME retains geometric and parametric information, PD-PME constructs the latent space purely from physical descriptors, which can be advantageous when physical trends are more relevant than shape variance. These distinctions highlight the versatility of PME-based approaches in design-space reduction tasks where physical interpretability and parameter reconstruction are essential.
@@
The effectiveness of PI-PME and PD-PME becomes particularly apparent in real-world applications such as bio-inspired underwater gliders, naval propellers, or classical airfoils (e.g., the RAE-2822). In these cases, fluid-dynamic effects strongly impact performance, and omitting them from the dimensionality reduction step may lead to suboptimal designs and increased resource expenditure. By embedding physical information—whether partially, as in PI-PME, or exclusively, as in PD-PME—into the reduced space, it becomes possible to filter the vast design domain more intelligently, ensuring that only the most relevant configurations are explored. This synergy between geometric and physical factors not only improves the quality of early design-stage decisions but also streamlines the path to final optimization, cutting down both computational time and overall development costs. It is important to emphasize that the test cases discussed throughout the paper are not simplified academic benchmarks, but rather representative of real-world early-stage design workflows in aerospace and marine engineering. Each case reflects a high-dimensional design problem where physical performance plays a central role: the RAE-2822 airfoil is a canonical transonic shape studied for aerodynamic efficiency; the seakeeping hull and propeller designs involve complex hydrodynamic effects under realistic operating conditions; and the glider case addresses the viscous performance of bioinspired geometries. In all scenarios, dimensionality reduction is not an optional step, but a necessary enabler to manage complexity, improve interpretability, and make downstream optimization tractable. PME’s flexible structure also makes it directly applicable to structural optimization. Instead of relying on aerodynamic or hydrodynamic simulations, one may incorporate structural properties (e.g., mass moments of inertia, stiffness distributions, or stress fields) into the PME framework, enabling integrated assessments of mechanical performance [22, 34].
@@
The following sections illustrate the mathematical underpinnings of PME, PI-PME, and PD-PME, including detailed discussions of how generalized PCA is adapted to incorporate physical data, and how these frameworks can be applied to a range of high-dimensional vehicle design problems. Through examples, the benefits of early-stage integration of physical insights will be demonstrated, showing improved design robustness, a decrease in exhaustive computational campaigns, and the potential for more agile innovation cycles in modern vehicle development. It may finally be noted that, while this work focuses on vehicle applications, the proposed method is broadly applicable to general shape optimization problems involving functional surfaces influenced by physical constraints.
@@ 2 Design-space dimensionality reduction methods
In shape optimization problems aimed at minimizing an objective function $f({\mathbf{u}})$, the optimum shape is uncertain before and during the optimization process. The design variable vector ${\mathbf{u}}\in\mathcal{U}\subset\mathbb{R}^{M}$, with $\mathcal{U}$ the design domain, can therefore be treated as a random variable, with an associated probability density $p({\mathbf{u}})$ which reflects the likelihood of a certain design being optimal, given the existing knowledge. If nothing is known about the design, $p$ can be uniform. This uncertainty propagates through the modeling process, generating uncertainties in both geometric configurations and the associated physical responses.
Dimensionality reduction techniques such as PCA, SVD, POD, and KLE provide a systematic way to extract the dominant directions of variability from such data, i.e., the combinations of design variables which most influence the results. These methods enable the construction of compact, low-dimensional representations that preserve the most informative features for design analysis and decision-making.
As will be discussed in the following sections, integrating physical insights into these dimensionality reduction frameworks, for design-space assessment and subsequent shape optimization, provides a more robust path toward identifying optimal solutions under uncertainty.
@@ 2.1 Parametric model embedding
PME [32] is a design-space dimensionality reduction method that extends the standard PCA approach by incorporating both shape deformations and design variables into a generalized feature space. Specifically, PME solves a generalized eigenvalue problem, applying a generalized PCA to an augmented matrix that includes the discretized shape deformation vector ${\mathbf{d}}$ and the original design variables ${\mathbf{u}}$. This extension allows PME to directly map the reduced design space back to the original design variables without the need for reparameterization, which is typically required in standard PCA approaches. As a result, PME offers a more robust and practical method for maintaining the integrity of the original design features while facilitating effective shape optimization.
@@
where ${\boldsymbol{\delta}}({\boldsymbol{\xi}},{\mathbf{u}})\in\mathbb{R}^{n}$ is the resulting shape modification vector, defined by arbitrary shape parameterization or modification method (e.g., CAD parameterization, Bezier surfaces, free-form deformation, NURBS, etc.).
@@
$\mathbf{D}$ $=\left[\begin{array}[]{ccc}\hat{{\mathbf{d}}}_{1}^{(1)}&\dots&\hat{{\mathbf{d}}}_{1}^{(S)}\\ \vdots&\vdots&\vdots\\ \hat{{\mathbf{d}}}_{n}^{(1)}&\dots&\hat{{\mathbf{d}}}_{n}^{(S)}\\ \end{array}\right]$ (2)
$=\left[\begin{array}[]{ccc}\hat{d}_{1,\xi_{1}}({\mathbf{u}}_{1})&&\hat{d}_{1,\xi_{1}}({\mathbf{u}}_{S})\\ \vdots&&\vdots\\ \hat{d}_{L,\xi_{1}}({\mathbf{u}}_{1})&&\hat{d}_{L,\xi_{1}}({\mathbf{u}}_{S})\\ \vdots&\dots&\vdots\\ \hat{d}_{1,\xi_{n}}({\mathbf{u}}_{1})&&\hat{d}_{1,\xi_{n}}({\mathbf{u}}_{S})\\ \vdots&&\vdots\\ \hat{d}_{L,\xi_{n}}({\mathbf{u}}_{1})&&\hat{d}_{L,\xi_{n}}({\mathbf{u}}_{S})\\ \end{array}\right]$
@@
$\mathbf{P}=\left[\begin{array}[]{c}\mathbf{D}\\ \mathbf{U}\end{array}\right]$ (3)
with
$\mathbf{U}=\left[\begin{array}[]{ccc}\hat{{\mathbf{u}}}^{(1)}&\dots&\hat{{\mathbf{u}}}^{(1)}\\ \end{array}\right]=\left[\begin{array}[]{ccc}\hat{u}_{1,1}&&\hat{u}_{1,S}\\ \vdots&\cdots&\vdots\\ \hat{u}_{M,1}&&\hat{u}_{M,S}\\ \end{array}\right]$ (4)
@@
$\mathbf{G}=\left[\begin{array}[]{cc}\mathbf{G}_{\mathbf{d}}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}\\ \end{array}\right]$ (7)
and
${\mathbf{Z}}=\left[{{\mathbf{z}}}_{1}\,\,\,\dots\,\,\,{{\mathbf{z}}}_{S}\right]\qquad\mathrm{with}\qquad{{\mathbf{z}}}_{k}=\left[\begin{array}[]{c}\mathbf{q}_{k}\\ {\mathbf{v}}_{k}\end{array}\right]$ (8)
is the eigenvectors matrix, whereas $\boldsymbol{\Lambda}$ is the associated eigenvalues matrix. Here, $\mathbf{q}_{k}$ and ${\mathbf{v}}_{k}$ represent the eigenvector components associated to the shape modification ${\mathbf{d}}$ and design variable ${\mathbf{u}}$ vectors, respectively. The matrix $\mathbf{G}_{\mathbf{d}}=\mathrm{diag}\left(\mathbf{G}_{1},\dots,\mathbf{G}_{n}\right)$ is block diagonal and has dimensionality $\left[nL\times nL\right]$, with each $\left[L\times L\right]$ $k$-th block being a diagonal matrix itself
@@
$\mathbf{Z}^{\star}=\mathbf{Z}\mathbf{\Gamma}^{-1},$ (12)
where
$\mathbf{\Gamma}=\mathrm{diag}\left(\left[\sqrt{\mathbf{z}_{1}^{\mathsf{T}}\mathbf{G}\mathbf{W}\mathbf{z}_{1}},\dots,\sqrt{\mathbf{z}_{nL}^{\mathsf{T}}\mathbf{G}\mathbf{W}\mathbf{z}_{nL}}\right]\right).$ (13)
@@
The solutions $\lambda_{k}$ and the corresponding normalized eigenvectors $\mathbf{v}_{k}^{\star}$ (columns component of $\mathbf{Z}^{\star}$ associated to the original design variables, see Eq. 7) are used to construct the reduced dimensionality representation of the original parameterization by means of the $N$ reduced design variables $\mathbf{x}=[x_{1},\dots,x_{N}]^{\mathsf{T}}$. Defining the desired confidence level $l$, with $0<l\leq 1$, the number of reduced design variables $N$ is chosen such that
@@
Thus design parameters $\mathbf{u}$ are assigned zero weight in the inner product definition, but they remain structurally embedded in the generalized eigenvalue problem. This allows PME to retain a direct mapping from the reduced space back to the original parametric domain, thus enabling analytical backmapping (see Eq. 15). As a result, the dimension reduction does not compromise the ability to reconstruct feasible and interpretable design configurations.
Finally, to ensure that all the samples in $\mathbf{P}$ can be reconstructed through the reduced-order representation of the original design space, the reduced design variables ${\mathbf{x}}$ can be bounded by identifying the minimum and maximum values attained by each component $\Theta_{jk}$ of the projection coefficients $\boldsymbol{\theta}_{j}$, as follows: